OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 9? is it x^2+y^2=81?

OpenStudy (lgbasallote):

\[\implies (x-0)^2 + (y-0)^2 = 9^2\implies x^2 + y^2 = 81\]

OpenStudy (anonymous):

so its rihgt ya?

OpenStudy (lgbasallote):

lol yeah